Angela plays soccer and golf for a total of 125 minutes every day. She plays soccer for 45 minutes longer than she plays golf.
Part A: Write a pair of linear equations to show the relationship between the number of minutes Angela plays soccer (x) and the number of minutes she plays golf (y) every day. (5 points)
Part B: How much time does Angela spend playing golf every day? Show your work. (3 points)
Part C: Is it possible for Angela to have spent 80 minutes playing soccer if she plays for a total of exactly 125 minutes and plays soccer for 45 minutes longer than she plays golf? Explain your reasoning. (2 points)

I tried to do it but then my work never made an sense

Answer by ikleyn(52775)

.
Angela plays soccer and golf for a total of 125 minutes every day. She plays soccer for 45 minutes longer than she plays golf.
Part A: Write a pair of linear equations to show the relationship between the number of minutes Angela plays soccer (x)
and the number of minutes she plays golf (y) every day. (5 points)
Part B: How much time does Angela spend playing golf every day? Show your work. (3 points)
Part C: Is it possible for Angela to have spent 80 minutes playing soccer if she plays for a total of exactly 125 minutes
and plays soccer for 45 minutes longer than she plays golf? Explain your reasoning. (2 points)
I tried to do it but then my work never made an sense
~~~~~~~~~~~~~~~~~

Part A)

Write equations as you read the problem


    x + y = 125   minutes    (1)    ("Angela plays soccer and golf for a total of 125 minutes every day.")
   
    x - y =  45   minutes    (2)    ("She plays soccer for 45 minutes longer than she plays golf.")


Thus equations are written and this part is DONE.


Part B)

To answer this question, you need to solve this system of equations for y.


It can be done in different ways.

One way is to use Elimination method.  For it, subtract equation (2) from equation (1).  You will get


    y - (-y) = 125 - 45;   or  2y = 80;   or  y = 80/2 = 40 minutes playing golf.


Another way is to use the Substitution method.


Part C)

To answer this question, you need to solve this system of equations for x.


It can be done in different ways.

One way is to use Elimination method.  For it, add equation (1) and (2).  You will get


    x + x = 125 + 45;   or  2x = 170;   or  x = 170/2 = 85 minutes playing soccer.


Another way is to use the Substitution method.


///////////////


Solved and explained.

If you want to see many other similar (and different) solved problems of this type. look into the lesson
    - Word problems that lead to a simple system of two equations in two unknowns
in this site.


Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic "Systems of two linear equations in two unknowns".

Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.